Fern has preferred stock selling for 93.3 percent of par that pays an 9.7 percent annual coupon (dividend rate). Assume the Par Value of the Preferred stock is $100. What would be Fern's component cost of preferred stock?
Answer #1

Dividend =9.7%*100 =9.7
Price =Par value*93.3% =100*93.30% =93.30
Cost of Preferred Stock =Dividend /Price =9.7/(93.3%*100) =10.40%